Use the following demand and supply functions to answer this question; Qd=100-2p; Qs=60+2p ; If a ceiling price of $8 is established by the government, then

a new equilibrium price will emerge.

a shortage of 84 will result.

a shortage of 8 will result.

a surplus of 8 will result.
Answer #1

IF the price ceiling is $8 the the demand will be 100 - 2(8) = 100 -16 = 84. And the supply will be 60 + 2(8) = 60 + 16 = 76.

Demand is 84 and the supply is 76. As the demand is more there will be a shortage of 8 will result. The answer is "C", A shortage of 8 will result.